Someone explain how to find x-intercept and y-intercept for equation 3x+y=2. Thanks
Answer by Cromlix(4381)

Sort the equation out to y = mx + c
3x + y = 2
y = -3x + 2
When the line cuts the y axis x=0
y = -3(0) + 2
y = 2 (0,2)
When the line cuts the x axis y=0
0 = -3x + 2
Rearrange
3x = 2
x = 2/3 (2/3,0)
Hope this helps.
:-)